# A first look at License Variants in the PyPI Ecosystem

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
@@ 1. Introduction
Even small changes in license text can introduce significant differences in terms and conditions. For instance, as shown in Figure 1, the Python package usd-core (usd, 2025b) (with over 580k downloads on PyPI (PyP, 2025c) and 1.3k forks on GitHub (usd, 2025a)) features a license that differs from the original Apache License 2.0 specifically in Trademarks, imposing more restrictive conditions, and overlooking this difference may lead to non-compliance of trademark using. On the other hand, some software owners, particularly commercial entities, choose not to use standard licenses and instead opt for completely custom licenses that better serve their interests (Meloca et al., 2018; Xu et al., 2023a; Lindman et al., 2010). For example, the cuda-python package (cud, 2025), a comprehensive toolkit providing Python access to the NVIDIA CUDA platform, uses a completely customized ”NVIDIA SOFTWARE LICENSE” that imposes significant restrictions, prohibiting modification and redistribution. Failure to accurately identify these terms could lead to serious consequences.
@@
Inspired by our findings, we propose LV-Parser, a license parsing method that explicitly accounts for license variants. By leveraging diff-based analysis, knowledge of standard licenses, and carefully crafted prompts for advanced language models, LV-Parser achieves both high efficiency and accuracy in handling diverse license texts. Building on LV-Parser’s capabilities, we further introduce LV-Compat, a more precise pipeline for detecting license incompatibilities in package dependency networks. Our evaluation demonstrates that LV-Parser achieves high accuracy (0.936) while reducing LLM query costs by nearly 30% compared to baseline methods, LV-Compat significantly improves incompatibility detection by identifying 5.2 times more incompatible packages than existing methods with 0.98 precision in its detection results.
@@ 2.1.1. License Modeling
However, this modeling approach has limitations in capturing fine-grained differences across licenses. For example, it cannot effectively distinguish between GPL and LGPL, particularly regarding their different treatments of dynamic linking. It also does not capture the stricter requirements of the AGPL, such as the obligation to disclose source code when software is accessed over a network. Additionally, the model fails to represent other critical legal aspects, such as the termination of patent grants upon litigation, which are present in licenses like Apache 2.0 and MPL 2.0. These omissions limit the model’s ability to support precise and reliable compatibility assessments.
To address these limitations, the studies (Xu et al., 2023d, c) proposed a more fine-grained license modeling approach. For example, the copyleft dimension (i.e., the requirement to apply the same license to derivative works) is categorized into permissive, file-level, library-level, and strong copyleft types, which is important in the context of package dependencies. Additional dimensions such as network-based source disclosure, termination of patent grants upon litigation, and enhanced attribution requirements are also incorporated. Therefore, in this work, we adopt and refine this license modeling approach. The detailed implementation is described in Section 4.
@@ 2.1.2. License Incompatibility Definition
In the context of package dependencies, Xu et al. (Xu et al., 2023d, c) introduces a more precise framework with two types of license compatibility: secondary compatibility and combinative compatibility. The definitions provided by previous works (Xu et al., 2023b, a; Cui et al., 2025; Liu et al., 2024b) correspond to the notion of secondary compatibility, where all parts of a derivative work can be re-licensed under the downstream license. By contrast, combinative compatibility allows for scenarios where a project as a whole may adopt a downstream license, while third-party components retain their original licenses and constraints. For example, an MIT-licensed project depending on an LGPL-licensed package fits the combinative compatibility category, as the LGPL component retains its original terms even when incorporated into the broader project. In this paper, we adopt this definition of compatibility, and the corresponding determination method will be described in Section 5.
@@ 3.1. Research question
    Rationale. Understanding how common license variants are, how they are distributed, and what textual differences they exhibit from standard licenses offers crucial empirical insights for designing more effective license analysis methods.
@@ 3.4.1. RQ1
Table 3. Compatibility-related part of License Model used in LV-Parser.
@@ 3.4.2. RQ2
As described in Section 2.1.2, the definition of secondary and combinative license compatibility fits well within the context of packaging ecosystems (Xu et al., 2023c), where a package may be considered a derivative work of its dependencies, according to the interpretation of the Free Software Foundation (FSF) (Wha, 2025). Based on this definition and the compatibility check methodology proposed by Xu et al. (Xu et al., 2023d), we manually review and annotate the 54 license texts along key legal dimensions, including copyright grant, copyleft level, patent grant, trademark grant, usage restrictions, etc. Note that our goal is to make a conservative estimate of incompatibilities within the PyPI ecosystem, so it is acceptable that the dimension set may not be strictly exhaustive. Then, we adopt widely accepted principles derived from previous work and official interpretations (Xu et al., 2023c, d; gpl, 2025) (e.g., FSF explicitly states that free software cannot depend on proprietary software due to fundamental license incompatibility (FSW, 2025)) to identify license incompatibilities of the 54 packages and their downstream releases. Any cases in which compatibility could not be determined with confidence are conservatively labeled as unknown.
Our analysis reveals that 234 (10.7%) out of the 2,177 downstream packages have license incompatibilities with the upstream variants. At the release level, 3,070 (9.0%) out of 34,004 releases are found to be incompatible, while 7,426 releases cannot be conclusively assessed and are marked as unknown. Notably, among the 3,070 incompatible releases, 1,879 cases (5.5% of the total) involve license variants that impose proprietary prohibitions on copyright-related rights, such as use, modification, or distribution. In these cases, the downstream packages are released under free software licenses while depending on upstream proprietary software, which is explicitly considered incompatible according to FSF guidance (FSW, 2025). The 7,426 releases (21.8 % of the total) that could not be conclusively assessed using current methods further underscore the significant challenges in license compatibility analysis when variants are involved. This high proportion of “unknown” cases demonstrates the limitation of existing compatibility analysis frameworks when confronted with non-standard license texts, highlighting the need for more sophisticated approaches to handle the legal ambiguity introduced by license variants.
@@ 4.1. License Modeling
License modeling serves as the foundation for understanding licenses and analyzing compatibility. As discussed in Section 2.1.1, existing modeling approach(Xu et al., 2023b; tld, 2025; Xu et al., 2023a) in Table 1 demonstrates significant limitations that result in incomplete license representations. To address these shortcomings, we adopt the modeling framework proposed in (Xu et al., 2023d, c) and further refine it based on insights from our empirical study. As shown in Table 3, all key terms in our license modeling approach are directly related to license compatibility.
@@ 4.3.2. Experiment Setup
Recent research has utilized LLMs for license parsing and showed promising results (Cui et al., 2025; Ke et al., 2025; Kahol et al., 2025). However, these works adopt the license modeling framework from (Xu et al., 2023b) as shown in Table 1, which differs from our approach and makes it difficult to directly compare with LV-Parser. To provide a fair baseline, we implement a method same as their core methodology: the baseline model is provided with rich context, including term definitions and value definitions, along with the target license text, and is asked to determine the appropriate value for each term. To ensure a fair comparison, the baseline uses the same term and value definitions as LV-Parser and employs the same LLM, GPT-4.1. The only difference is that the baseline does not incorporate diff-based processing or provide similar example clauses as additional context. In all experiments, the LLM operates in greedy decoding mode with a temperature setting of zero. Additionally, we evaluate our approach by replacing GPT-4.1 with several different LLMs (including Claude-4-Sonnet (cla, 2025), DeepSeek-v3 (Liu et al., 2024a), Qwen3(Yang et al., 2025)) to demonstrate the robustness and generalizability of our parsing methodology, regardless of the specific model used.
@@ 5. The LV-Compat Pipeline
5.1. Incompatibility Determine Method
@@
Our algorithms, presented in Algorithm 1, center on copyleft terms as the primary determinants of compatibility in package dependencies. The compatibility check begins by extracting core terms from both licenses, with special focus on their copyleft levels. For exact license matches, both compatibility types are automatically satisfied. In other cases, the algorithm evaluates whether the downstream license can satisfy the upstream license’s conditions, considering secondary compatibility when all terms can be satisfied, and combinative compatibility for scenarios like MIT-licensed projects depending on LGPL components.
@@ 5.3. Evaluation of LV-Compat
To evaluate the effectiveness of LV-Compat in assessing the compliance of real-world software packages, we use the same 74 packages as described in Section 4.3. From all 3,449 of their downstream releases, we randomly sample 75 releases to serve as the evaluation set for LV-Compat. We compare LV-Compat with SILENSE(Xu et al., 2023c), which utilizes the same license modeling approach and license definitions. As described in Section3.3.2, SILENSE obtains license information without considering variants. Furthermore, its compatibility analysis is limited to the predefined set of 63 licenses, meaning that any licenses outside this set cannot be assessed for compatibility. The results, shown in Table 5, highlight that LV-Compat substantially outperforms SILENSE in real-world compliance assessment. Among the 75 downstream packages evaluated, each package was checked for compatibility with all its dependencies. “Package” columns in Table 5 reflect cases where packages were found to have incompatible, unknown, or compatible dependencies.
LV-Compat identified 52 packages with at least one incompatible dependency, compared to just 10 for SILENSE. In addition, SILENSE was unable to determine compatibility for 65 packages, whereas LV-Compat reduced the number of packages marked as “Unknown” to only 7. On the dependency level, the “Dependency” columns in the table record the total number of dependency relationships found incompatible or unknown. LV-Compat detected 116 incompatible dependencies, far more than the 30 identified by SILENSE, and decreased the number of dependencies marked ”Unknown” from 598 to 52. A key advantage of LV-Compat is its ability to parse and analyze third-party software licenses embedded within dependency license text, a dimension SILENSE does not cover. This further analysis uncovered 238 incompatibilities and 148 unknowns relating to third-party components. By incorporating third-party software considerations, LV-Compat delivers a far more comprehensive assessment of license compatibility and compliance compared to prior approaches.
@@
Incom. Unknown Compat. Incom. Unknown Incompat. Unknown
@@ 6.1. Case study
There are also examples where the complexity of legal language or the need for deeper legal expertise leads to disagreements even among authors during annotation. For instance, in the license of the package encodec(enc, 2025), the clause ”Patent and trademark rights are not licensed under this Public License.” was annotated by human annotators as ”not mentioned” for the patent grant term, while LV-Parser classified it as ”explicitly not granted.” According to (Meeker, 2020), even in the absence of an explicit patent clause, most licenses are considered to implicitly grant some patent rights, although there is little relevant case law. Therefore, after discussion, the annotators chose to mark the clause as 0 (”not mentioned”) rather than -1 (”explicitly not granted”).

### v2

# Small Changes, Big Trouble: Demystifying and Parsing License Variants for Incompatibility Detection in the PyPI Ecosystem

@@ Abstract.
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: 2026 IEEE/ACM 48th International Conference on Software Engineering; April 12–18, 2026; Rio de Janeiro, Brazilisbn: 978-1-4503-XXXX-X/2018/06
@@ 1. Introduction
Even small changes in license text can introduce significant differences in terms and conditions. For instance, as shown in Figure 1, the Python package usd-core (usd, 2025b) (with over 580k downloads on PyPI (PyP, 2025c) and 1.3k forks on GitHub (usd, 2025a)) features a license that differs from the original Apache License 2.0 specifically in Trademarks, imposing more restrictive conditions, and overlooking this difference may result in non-compliance with trademark usage requirements. On the other hand, some software owners, particularly commercial entities, choose not to use standard licenses and instead opt for completely custom licenses that better serve their interests (Meloca et al., 2018; Xu et al., 2023a; Lindman et al., 2010). For example, the cuda-python package (cud, 2025), a comprehensive toolkit providing Python access to the NVIDIA CUDA platform, uses a completely customized ”NVIDIA SOFTWARE LICENSE” that imposes significant restrictions, prohibiting modification and redistribution. Failure to accurately identify these terms could lead to serious consequences.
@@
Inspired by our findings, we propose LV-Parser, a license parsing method that explicitly accounts for license variants. By leveraging diff-based analysis, knowledge of standard licenses, and carefully crafted prompts for advanced language models, LV-Parser achieves both high efficiency and accuracy in handling diverse license texts. Building on LV-Parser’s capabilities, we further introduce LV-Compat, a more precise pipeline for detecting license incompatibilities in package dependency networks. Our evaluation demonstrates that LV-Parser achieves high accuracy (0.936) while reducing LLM query costs by nearly 30% compared to baseline methods. LV-Compat significantly improves incompatibility detection by identifying 5.2 times more incompatible packages than existing methods with 0.98 precision in its detection results.
@@ 2.1.1. License Modeling
However, this modeling approach has limitations in capturing fine-grained differences across licenses. For example, it cannot effectively distinguish between GPL and LGPL, particularly regarding their different treatments of dynamic linking. It also does not capture the stricter requirements of the AGPL, such as the obligation to disclose source code when software is accessed over a network. Additionally, the model fails to represent other critical legal aspects, such as the termination of patent grants upon litigation, which are present in licenses such as Apache 2.0 and MPL 2.0. These omissions limit the model’s ability to support precise and reliable compatibility assessments.
To address these limitations, Xu et al. (Xu et al., 2023d, c) propose a more fine-grained license modeling approach. For example, the copyleft dimension (i.e., the requirement to apply the same license to derivative works) is categorized into permissive, file-level, library-level, and strong copyleft types, which is important in the context of package dependencies. Additional dimensions such as network-based source disclosure, termination of patent grants upon litigation, and enhanced attribution requirements are also incorporated. Therefore, in this work, we adopt and refine this license modeling approach. The detailed implementation is described in Section 4.
@@ 2.1.2. License Incompatibility Definition
In the context of package dependencies, Xu et al. (Xu et al., 2023d, c) introduce a more precise framework with two types of license compatibility: secondary compatibility and combinative compatibility. The definitions provided by previous works (Xu et al., 2023b, a; Cui et al., 2025; Liu et al., 2024b) correspond to the notion of secondary compatibility, where all parts of a derivative work can be re-licensed under the downstream license. By contrast, combinative compatibility allows for scenarios where a project as a whole may adopt a downstream license, while third-party components retain their original licenses and constraints. For example, an MIT-licensed project depending on an LGPL-licensed package fits the combinative compatibility category, as the LGPL component retains its original terms even when incorporated into the broader project. In this paper, we adopt this definition of compatibility, and the corresponding determination method is described in Section 5.
@@ 3.1. Research question
    Rationale. Understanding the prevalence and distribution of license variants, as well as their textual deviations from standard licenses, provides crucial empirical insights for designing more effective license analysis methods.
@@ 3.4.1. RQ1
Table 3. License Model used in LV-Parser.
@@
Definition Clause that provides the definition, meaning, or clarification of a specific term, concept, or entity used within the license. 0:Not mentioned or missing definition clause. 1:Explicit definition clause exists.
@@
Termination for Breach A clause specifying that the license and associated rights terminate if the licensee breaches the license terms. 0: No explicit termination-for-breach. 1: Explicit license termination clause for breach.
@@
Disclaimer A clause disclaiming warranties or liabilities, stating that the work is provided ’as is’, and that the licensors make no representations or warranties with respect to the work. 0: No warranty disclaimer. 1: Explicit no warranty/liability clause.
Governing Law A clause specifying which country’s or region’s law applies to the license or to disputes arising under it. None or a list: Explicitly specifies governing law.
Instruction A clause describing instructions or guidance on how to use this license. 0: No explicit instruction/guidance. 1: Explicitly gives instructions/guidance.
@@
Other Any other provisions not covered by the above categories. -
@@ 3.4.2. RQ2
As described in Section 2.1.2, the definition of secondary and combinative license compatibility fits well within the context of packaging ecosystems (Xu et al., 2023c), where a package may be considered a derivative work of its dependencies, according to the interpretation of the Free Software Foundation (FSF) (Wha, 2025). Based on this definition and the compatibility check methodology proposed by Xu et al. (Xu et al., 2023d), we manually review and annotate the 54 license texts along key legal dimensions, including copyright grant, copyleft level, patent grant, trademark grant, usage restrictions, etc. Note that our goal is to make a conservative estimate of incompatibilities within the PyPI ecosystem, so it is acceptable that the dimension set may not be strictly exhaustive. Then, we adopt widely accepted principles derived from previous work and official interpretations (Xu et al., 2023c, d; gnu, 2025a; apa, 2025; cec, 2025; gnu, 2025b; ecl, 2025; eup, 2025) (e.g., FSF explicitly states that free software cannot depend on proprietary software due to fundamental license incompatibility (gnu, 2025a)) to identify license incompatibilities of the 54 packages and their downstream releases. Any cases in which compatibility could not be determined with confidence are conservatively labeled as unknown.
Our analysis reveals that 234 (10.7%) out of the 2,177 downstream packages exhibit license incompatibilities with their upstream variants. At the release level, 3,070 (9.0%) out of 34,004 releases are found to be incompatible, while 7,426 releases cannot be conclusively assessed and are labeled as “unknown”. Among the 3,070 incompatible releases, 1,879 cases (5.5% of the total) involve license variants that impose proprietary prohibitions on copyright-related rights such as use, modification, or distribution. In these cases, downstream packages are released under free software licenses while depending on upstream proprietary software—explicitly deemed incompatible under FSF guidance (gnu, 2025a).
While standard SPDX licenses can themselves lead to downstream incompatibilities, modifications in variants greatly amplify this complexity. In our analysis of licenses altered relative to their SPDX originals, we identify two contrasting patterns: some variants introduce stricter obligations, leading to a 21% increase in incompatible downstream releases compared with their SPDX counterparts, whereas others add exceptions that relax obligations, resulting in a 76% reduction in incompatibilities. These findings indicate that license variants can exert opposite effects on compatibility and underscore the need for dedicated analysis methods that can effectively capture such modifications to improve recall while reducing false positives. The 7,426 uncertain releases (21.8% of the total) further highlight the substantial challenges and the limitations of current compatibility frameworks when confronted with non-standard or modified license texts.
@@ 4.1. License Modeling
License modeling serves as the foundation for understanding licenses and analyzing compatibility. As discussed in Section 2.1.1, existing modeling approach (Xu et al., 2023b; tld, 2025; Xu et al., 2023a) in Table 1 demonstrates significant limitations that result in incomplete license representations. To address these shortcomings, we adopt the modeling framework proposed in (Xu et al., 2023d, c) and further refine it based on insights from our empirical study. As shown in Table 3, it includes both compatibility-related and unrelated terms, providing a complete overview of the terminology used in our license modeling.
@@ 4.3.2. Experiment Setup
Recent research has utilized LLMs for license parsing and showed promising results (Cui et al., 2025; Ke et al., 2025; Kahol et al., 2025). However, these works adopt the license modeling framework from  (Xu et al., 2023b) as shown in Table 1, which differs from our approach and makes it difficult to directly compare with LV-Parser. To provide a fair baseline, we adopt their core methodology as the baseline implementation: the baseline model is provided with rich context, including term definitions and value definitions, along with the target license text, and is asked to determine the appropriate value for each term. To ensure a fair comparison, the baseline uses the same term and value definitions as LV-Parser and employs the same LLM, GPT-4.1. The only difference is that the baseline does not incorporate diff-based processing or provide similar example clauses as additional context. In all experiments, the LLM operates in greedy decoding mode with a temperature setting of zero. Additionally, we evaluate our approach by replacing GPT-4.1 with several different LLMs (including Claude-4-Sonnet (cla, 2025), DeepSeek-v3 (Liu et al., 2024a), Qwen3 (Yang et al., 2025)) to demonstrate the robustness and generalizability of our parsing methodology, regardless of the specific model used.
@@ 5. The LV-Compat Pipeline
5.1. Incompatibility Determination Method
@@
Our algorithms, presented in Algorithm 1, center on copyleft terms as the primary determinants of compatibility in package dependencies. The compatibility check begins by extracting core terms from both licenses, with special focus on their copyleft levels. For exact license matches, both compatibility types are automatically satisfied. In other cases, the algorithm evaluates whether the downstream license can satisfy the upstream license’s conditions, considering secondary compatibility when all terms can be satisfied, and combinative compatibility for scenarios such as MIT-licensed projects depending on LGPL components.
@@ 5.3. Evaluation of LV-Compat
To evaluate the effectiveness of LV-Compat in assessing the compliance of real-world software packages, we use the same 74 packages as described in Section 4.3. From all 3,449 of their downstream releases, we randomly sample 75 releases to serve as the evaluation set for LV-Compat. We compare LV-Compat with SILENCE (Xu et al., 2023c), which utilizes the same license modeling approach and license definitions. As described in Section 3.3.2, SILENCE obtains license information without considering variants. Furthermore, its compatibility analysis is limited to the predefined set of 63 licenses, meaning that any licenses outside this set cannot be assessed for compatibility. The results, shown in Table 5, highlight that LV-Compat substantially outperforms SILENCE in real-world compliance assessment. Among the 75 downstream packages evaluated, each package was checked for compatibility with all its dependencies. “Package” columns in Table 5 reflect cases where packages were found to have incompatible, unknown, or compatible dependencies.
LV-Compat identified 52 packages with at least one incompatible dependency, compared to just 10 for SILENCE. In addition, SILENCE was unable to determine compatibility for 65 packages, whereas LV-Compat reduced the number of packages marked as “Unknown” to only 7. On the dependency level, the “Dependency” columns in the table record the total number of dependency relationships found incompatible or unknown. LV-Compat detected 116 incompatible dependencies, far more than the 30 identified by SILENCE, and decreased the number of dependencies marked ”Unknown” from 598 to 52. A key advantage of LV-Compat is its ability to parse and analyze third-party software licenses embedded within dependency license text, a dimension SILENCE does not cover. This further analysis uncovered 238 incompatibilities and 148 unknowns relating to third-party components. By incorporating third-party software considerations, LV-Compat delivers a far more comprehensive assessment of license compatibility and compliance compared to prior approaches.
@@
Incom. Unknown Compat. Incom. Unknown Incom. Unknown
@@ 6.1. Case study
There are also examples where the complexity of legal language or the need for deeper legal expertise leads to disagreements even among authors during annotation. For instance, in the license of the package encodec (enc, 2025), the clause ”Patent and trademark rights are not licensed under this Public License.” was annotated by human annotators as ”not mentioned” for the patent grant term, while LV-Parser classified it as ”explicitly not granted.” According to Meeker (Meeker, 2020), even in the absence of an explicit patent clause, most licenses are considered to implicitly grant some patent rights, although there is little relevant case law. Therefore, after discussion, the annotators chose to mark the clause as 0 (”not mentioned”) rather than -1 (”explicitly not granted”).
@@ 7. Conclusion
8. Data Availability
We make our dataset and code available at https://figshare.com/s/4fbaedbeb120d1940f12.
Acknowledgements.
This work is sponsored by the National Natural Science Foundation of China 62332001, 62502030, and Fundamental Research Funds for the Central Universities FRF-TP-25-030.